### v1

# STEP: Simultaneous Tracking and Estimation of Pose for Animals and Humans

Abstract
We introduce STEP, a novel framework utilizing Transformer-based discriminative model prediction for simultaneous tracking and estimation of pose across diverse animal species and humans. We are inspired by the fact that the human brain exploits spatiotemporal continuity and performs concurrent localization and pose estimation despite the specialization of brain areas for form and motion processing. Traditional discriminative models typically require predefined target states for determining model weights, a challenge we address through Gaussian Map Soft Prediction (GMSP) and Offset Map Regression Adapter (OMRA) Modules. These modules remove the necessity of keypoint target states as input, streamlining the process. Our method starts with a known target state initialized through a pre-trained detector or manual initialization in the initial frame of a given video sequence. It then seamlessly tracks the target and estimates keypoints of anatomical importance as output for subsequent frames. Unlike prevalent top-down pose estimation methods, our approach doesn’t rely on per-frame target detections due to its tracking capability. This facilitates a significant advancement in inference efficiency and potential applications. We train and validate our approach on datasets encompassing diverse species. Our experiments demonstrate superior results compared to existing methods, opening doors to various applications, including but not limited to action recognition and behavioral analysis.
keywords:
Pose Estimation, Visual Tracking, Target Model Prediction.
1 Introduction
Pose estimation is a fundamental and demanding task in computer vision, with diverse practical applications in the real world. When applied to video sequences, pose estimation proves invaluable across various uses, including but not limited to action recognition, behavioral analysis, robotics, and surveillance, demonstrating its wide-ranging utility. Analyzing poses across videos involves tackling specific challenges, such as object detection, precise frame-to-frame tracking, and anatomical keypoint localization.
Motivation. Traditional top-down Pose Estimation Techniques require precise bounding box data, sourced either from automated detectors or manually annotated ground truths, to appropriately crop, center, and scale the images of the targets. These methods are designed with specific cropping standards that reflect established priors regarding the target shapes. Although these methods have been effective, their performance can diminish when encountering targets with non-standard aspect ratios, leading to less accurate pose estimation. To this end, we propose a transformer-based method to simultaneously handle tracking and pose estimation for the object within a video by concurrently estimating weights for three tasks: localization, tracking, and pose estimation. Integrating tracking eliminates the need for per-frame object detections used in existing top-down methods. Unlike these conventional approaches, our proposed method offers a robust alternative not hindered by variations in aspect ratios, enhancing its utility across diverse imaging conditions and target configurations.
Pose Estimation. Pose estimation approaches can be categorized into two main paradigms: top-down and bottom-up. The top-down approaches [47] [38] first detect the object of interest employing object detectors and subsequently conduct pose estimation for each detection. In contrast, the bottom-up approaches [34] [51] [16] directly regress the keypoints and later concentrate on associating the detected keypoints to the object. The top-down approach achieves accuracy at a higher cost with an extra detection step, while the bottom-up paradigm, though more efficient, attains lower accuracy. Traditional top-down Pose Estimation Techniques require precise bounding box data typically, a 1.33 height-to-width ratio for human subjects to accommodate their vertical orientation and a 1.33 width-to-height ratio for animals to capture their broader horizontal spread. CNN-based models used for regressing keypoints in pose estimation, like [39] [42] [44], face challenges in capturing image-specific dependencies because of static convolutional kernels. Furthermore, these models excel in revealing class-specific saliency maps rather than at establishing relationships between structural elements, such as the connection between anatomical structure and keypoint locations within the pose context [49].
Visual Object Tracking. Deep neural networks have shown remarkable performance in object tracking and pose estimation tasks. Within the tracking domain, two predominant methodologies have emerged: Siamese networks [22] and networks integrating discriminative appearance modules [2]. Object Tracking deals with accurately identifying and tracking each target object within video streams. This field has progressed significantly due to advances in object detection technologies, with most modern trackers adopting the tracking-by-detection approach. This technique applies an object detector across successive frames to generate detections that are then used for tracking. Centernet [13] is commonly employed in many tracking-by-detection systems due to its operational efficiency, whereas YOLOX [14] has gained popularity for its exceptional detection accuracy. Typically, these methods incorporate the Kalman filter [43] as a motion model to predict future tracklet positions. The Intersection over Union (IoU) metric is used to gauge the similarity between predicted and detected boxes for a reliable association. However, the Kalman filter is noted for its vulnerabilities, such as susceptibility to state noise and amplification of temporal discrepancies. Furthermore, the application of the Kalman filter, despite its widespread use, requires careful tuning of its parameters to mitigate the effects of its inherent vulnerabilities. To address these issues, some tracking systems have begun integrating more sophisticated deep learning-based motion models, which can offer improved robustness and accuracy in dynamic environments.
Contribution. Existing top-down approaches utilize object detectors like [18] [28] to determine a bounding box around the object for pose estimation. Object detectors primarily focus on recognizing and localizing objects within individual images rather than maintaining a (tracking) identity across frames, creating a gap in pose analysis within video sequences. Our proposed approach addresses this gap through tracking capability while simultaneously estimating pose keypoints. Our approach uses a transformer-based method to simultaneously handle tracking and pose estimation for the object within a video by concurrently estimating weights for three tasks: localization, tracking, and pose estimation. Integrating tracking eliminates the need for per-frame object detections used in existing top-down methods. In summary, our contributions are as given below.
  • We conduct a comprehensive study on the pose estimation task within a discriminative model prediction framework. We utilize a Transformer-based model prediction module while simultaneously tracking the object of interest— an unexplored area in previous research.
  • We introduce the Gaussian Map Soft Prediction (GMSP) and Offset Map Regression Adapter (OMRA) modules, which work together to accurately localize key points corresponding to the target state, ensuring both efficiency and precision in pose key point estimation.
  • In the inference phase, we propose a confidence-based memory update of our tracker. This update integrates previous frame outputs to supervise and estimate subsequent outputs effectively. We showcase the effectiveness of our approach through a case study on Awaji Monkey Center (AMC) video streams.
We begin by reviewing relevant works on tracking and pose estimation closely related to ours in Section 2. Following that, Section 3 provides a detailed explanation of our proposed methodology, including the network architecture and the steps involved in training STEP and performing inference. We demonstrate that our approach delivers competitive performance in pose estimation compared to state-of-the-art methods while also incorporating tracking. Notably, our method surpasses these approaches in top-down evaluations. Finally, Section 4 presents experiments and exhaustive ablation studies to validate our findings further.
2 Related Works
Our proposed approach is inspired by established discriminative model prediction paradigms used in visual object tracking [2] [8] [7] [31]. We explore the potential of this framework to predict keypoints for pose estimation while performing tracking simultaneously. In contrast to prevailing trends, which often focus on solving dedicated tasks sequentially, we demonstrate how vision transformer-based discriminative model weight predictions bridge the gap, enabling simultaneous predictions instead of sequential estimations. Vision Transformers have demonstrated significant promise across various other related vision tasks, such as Recognition [12], Object Detection [4], and Segmentation [36]. Recently, numerous studies have explored the potential of Vision Transformers for tracking [5] and Pose estimation [47] [49]. The discriminative features predicted by transformers for bounding box regression in tracking have demonstrated more promising outcomes than predictions from traditional optimization-based discriminative models. Additionally, in pose estimation, the transformer architecture [40] holds a distinct advantage over CNNs due to its attention layers, facilitating the model in capturing interactions among any pair of keypoint locations. We now present an overview of existing literature emphasizing Transformers within the domains of pose estimation and tracking tasks on video sequences.
2.1 Vision Transformers for Tracking
Modern trackers such as [4] [5] [52] utilize transformers to predict discriminative features for target localization and bounding box regression. These trackers involve a transformer encoder that encodes target-related features, followed by a transformer decoder that combines training and test features using cross-attention layers to compute discriminative features [41] [52] [5]. Two specialized modules are employed by [52] for tasks of localization and bounding box regression, while [5] utilizes a feature fusion module. In [56], a Multi-object tracker (MOT) is introduced, drawing inspiration from tracking by detection. Additional works sharing a similar approach, such as [8] [2], incorporate a memory state to update online trackers. Expanding on the transformer architecture introduced in [4] [31] [32] presents a transformer-based model predictor that estimates weights for localization and bounding box regression tasks.
Figure 1: The proposed STEP framework for simultaneous tracking and pose estimation. (a) the full architecture, (b) the calculation of Target State Encodings, (c) the architecture of the offset map regression adapter (OMRA) module, and (e) the standard transformer architecture comprising multi-headed self-attention blocks.
Figure 2: Modules of STEP architecture are shown (a) GMSP Module, (b) Localizer Module $\Lambda_{L}$, (c) Keypoint Localizer Module $\Lambda_{K}$, and (d) Bounding Box Regressor Module $\Lambda_{B}$.
2.2 Vision Transformers for Pose Estimation
The evolution of pose estimation methods has swiftly transitioned from CNNs [44] to vision transformer-based networks due to their robust feature representation capabilities. In contrast to how transformers are used in discriminative model prediction-based tracking approaches, studies in pose estimation view transformers as more effective decoders [25] [27] [26]. In [49], the transformer architecture processes feature to capture global relationships within images. Introducing a token-based representation, [27] models relationships among pose keypoints. Additionally, HRFormer [54] utilizes transformers to learn high-resolution features. These transformer-based pose estimation methods outperform widely used keypoint estimation benchmarks compared to CNN-based methods. Investigating the capabilities of straightforward vision transformers, [47] explores simplified transformer baselines, eliminating heavyweight CNN-based feature extractors or intricate transformer architectures present in existing neural models. More recently, pose estimation in challenging scenarios, such as low-light images and limited annotations, has been addressed in [1] and [23].
2.3 Pose Estimation and Tracking in videos
Object tracking and pose estimation in videos have extensive applications, including action recognition [3] and the analysis of human or animal behavior in various environments. Notably, markerless pose estimation is gaining popularity in neuroscience and the study of animal behaviors [6]. Expanding on the widely used CNN-based architecture of the DeepLabCut model [30], [21] extends it to facilitate multi-animal tracking and pose estimation. SLEAP [33] employs a lightweight CNN, effectively tracking insects (drosophila, bees) and rodents. In these approaches, pose estimation occurs frame-by-frame, while the tracking framework relies on a computationally intensive optical flow-based method. An encoder-decoder setup is proposed by [37] to encode input frames and train the model for reconstructing spatiotemporal differences between video frames for keypoint discovery. Several studies have concentrated on analyzing human poses across videos. Approaches such as [44] [49] [54] involve framewise detections followed by dedicated modules for identification across frames.
3 Methodology
This work introduces STEP, a Transformer-based target model prediction network that conducts simultaneous tracking and pose estimation by predicting keypoints. We first provide a concise overview of existing Optimization and Transformer-based model predictors and then elaborate on our proposed approach, STEP, outlining its network architecture, training, and inference procedures in detail.
3.1 Background
Existing discriminative models like [2] [8] [7] [57] typically adhere to a shared foundational framework. They aim to solve optimization problems, as presented in Equation 1, where the target model $\mathcal{T}(.)$ generates the intended target states.
$w=\arg\min_{\hat{w}}\sum_{(x,y)\in\mathcal{S}_{train}}\phi(\mathcal{T}(\hat{w};x),y)+\lambda r(\hat{w})$ (1)
Given a collection of training samples comprising a total of $m$ training frames, denoted as $\mathcal{S}_{train}=\{(x_{i},y_{i})\}_{i=1}^{m}$, where $x_{i}\in\mathcal{X}$ represents the high-dimensional feature map of the $i$-th frame and $y_{i}\in\mathcal{Y}$ represents the corresponding desired target state. The optimization problem outlined in Equation 1 aims to attain the optimal weight $w$ for the target model $\mathcal{T}$. This is achieved by minimizing residuals computed by the function $\phi(.)$, which measures the error between the output of the target model and the ground truth labels $y$. $r(.)$ is a regularization term weighted by a scalar $\lambda$. Approaches following this optimization strategy optimize weights with limited information, relying solely on the provided training frames. Furthermore, they struggle to incorporate learned priors into the target model. Rather than directly minimizing the objective outlined in Equation 1, [31] introduces a Transformer-based model predictor approach. This method predicts target model weights solely from data through end-to-end learning. Nevertheless, their primary emphasis lies in learning features that effectively differentiate the target state from previously encountered states within $\mathcal{S}_{train}$. For pose estimation, the model predictor should precisely capture crucial information relevant to anatomical keypoints rather than solely concentrating on localization tasks.
Figure 3: For a video sequence, columns display target score maps $\hat{K}_{\text{gm}}$ when initiated with bounding boxes, as depicted in (a) and (b). In (c), our STEP framework runs concurrently for each bounding box to track and estimate the pose of respective targets. Notably, observe the robustness of STEP in estimating keypoints, particularly in scenarios involving occlusion. A failure case is showcased in the last frame of (c) for the red bounded target, stemming from ambiguous activations within the red target’s score map, as evident in the last frame of (a). Additional qualitative results across various species are provided in the Supplementary material.
3.2 Network Architecture
Our proposed network adopts an end-to-end learning approach to predict the weights of the target model purely from data. The overview of our methodology is depicted in Figure 1. We leverage a baseline network pre-trained on Imagenet [9] to extract high-dimensional image features. We closely adhere to the discriminative model prediction approach introduced by [31] and expand it to learn target model weights specifically for the keypoint regression task. Our rationale is rooted in the belief that this discriminative model prediction approach captures more target-specific information, making it suitable for distinguishing keypoints of anatomical importance. Unlike [31], our target state information includes information on $k$-keypoints crucial for the pose estimation task. The main components of our network are now described in detail.
3.2.1 Target State Encodings
Adhering to the established principles outlined in [46] for precise tracking, tracking methodologies typically divide the task into two sub-tasks: localization and bounding box regression. Building upon this widely adopted approach, it becomes imperative to incorporate an additional step in our simultaneous pose estimation methodology, which involves deriving regression targets for estimating keypoint locations. Consider $K^{t}\in\mathrm{R}^{2\times k}$ as the ground truth representation of keypoints for the target object, comprising $k$ keypoints, each denoted by $K_{i}^{t}=(x_{i},y_{i})$. Here, $(x_{i},y_{i})$ represents the coordinates of the $i$-th keypoint. We use the subscript $t$ to highlight that the keypoints belong to the target object among the multiple object annotations present in the frames. Given a training frame $I\in\mathrm{R}^{H_{i}\times W_{i}\times 3}$, the pre-trained backbone network extracts its features $X\in\mathrm{R}^{H_{f}\times W_{f}\times n}$ in $n$-dimensional space, scaling down the image space by factor $s$. We represent $K_{i}^{t}$ at each pixel location $(j^{x},j^{y})$ on the feature map $X$ as follows,
$\kappa_{i}=\left(\left\lfloor\dfrac{s}{2}\right\rfloor+s\cdot j^{x}-x_{i}),\left\lfloor\dfrac{s}{2}\right\rfloor+s\cdot j^{y}-y_{i})\right)$ (2)
where $H_{i}=s\cdot H_{f}$ and $W_{i}=s\cdot W_{f}$. This representation of keypoint $K^{t}$ is denoted as offset map $\mathrm{K}^{t}_{\text{om}}=(\kappa_{i})_{i=1}^{k}$ with $\mathrm{K}^{t}_{\text{om}}\in\mathrm{R}^{H_{f}\times W_{f}\times 2k}$. Additionally, apart from the keypoint offset map $\mathrm{K}^{t}_{\text{om}}$ of the target, we generate a $k$-channel Gaussian map $\mathrm{K}_{\text{gmsp}}\in\mathrm{R}^{H_{f}\times W_{f}\times k}$, where each channel $i$ is a Gaussian map centered at the coordinate location of the $i$-th keypoint for all possible objects annotated in ground-truth. We create $K^{t}_{\text{gm}}$ by retaining the Gaussian map for keypoints related to the object of interest (target object) and set all other Gaussian maps to zero. Adopting the widely recognized LTRB representation used in [31] [52], a bounding box $\beta=\{b_{x_{1}},b_{y_{1}},b_{x_{2}},b_{y_{2}}\}$ in $xyxy$ format can be translated into the offset map $\mathrm{B}_{\text{om}}=\{\lfloor(s/2)\rfloor+s\cdot j^{x}-b\}\;\;\forall\;b\in\beta$. This $\mathrm{B}_{\text{om}}\in\mathrm{R}^{H_{f}\times W_{f}\times 4}$ is the offset map representation of the bounding box $\beta$. Moreover, we generate a Gaussian map $\mathrm{B}_{\text{gm}}$ centered at the center of the bounding box $\beta$, specifically for learning the localization task. In Figure 1(b), this estimation is illustrated by $\Pi_{b}$ and $\Pi_{kp}$ blocks. These maps, $\mathrm{K}^{t}_{\text{om}},\mathrm{K}^{t}_{\text{gm}},\mathrm{K}_{\text{gmsp}},\mathrm{B}_{\text{om}}$, and $\mathrm{B}_{\text{gm}}$ act as regression targets for the localizer and regression modules, described in Section 3.3.
We utilize two learnable embeddings in $n$ dimensions: $\phi_{\text{loc}}\in\mathrm{R}^{1\times n}$ and $\phi_{\text{kp}}\in\mathrm{R}^{k\times n}$. These embeddings are employed to distinguish between foreground and background information concerning target centers and keypoints, where $k$ represents the number of keypoints. For a given training frame $i$, we combine the extracted deep features $X$ with this target state information to create the training features for the Transformer module, as denoted in Equation 3.
$f_{i}=X+\Psi_{b}(B_{\text{om}})+\Psi_{\text{kp}}(K^{t}_{\text{om}})+\phi_{\text{loc}}\cdot B_{\text{gm}}+\phi_{kp}\cdot K^{t}_{\text{gm}}$ (3)
Here, $\Psi$ denotes a neural network containing multiple Multilayer Perceptrons (MLPs) designed to map the offsets into higher dimensions with $n$ channels, while the $\cdot$ operation signifies element-wise multiplication. Equation 3’s target states rely on $K_{\text{om}}$ and $K_{\text{gm}}$, necessitating initialization with keypoint data. We propose GMSP and OMRA Modules to eliminate this dependency (Section 3.2.3). Assuming our GMSP module infers enriched information approximating Gaussian maps $K_{\text{gmsp}}$, we adjust Equation 3 as follows:
$f_{i}=X+\Psi_{b}(B_{\text{om}})+\phi_{\text{loc}}\cdot B_{\text{gm}}+\phi_{kp}\cdot K_{\text{gmsp}}$ (4)
Moreover, we incorporate the test frame encoding, composed of deep features $X_{\text{test}}$ and a learnable token $\phi_{\text{test}}$, resulting in $f_{\text{test}}=X_{\text{test}}+\mu(\phi_{\text{test}})$. Here, $\phi_{\text{test}}$ represents the learnable token, and $\mu(\cdot)$ signifies the repetition of the token for each dimension of $X_{\text{test}}$.
3.2.2 Transformer Encoder and Decoder
The transformer encoder and decoder utilize training features $f_{i}$, and test features $f_{\text{test}}$ as inputs. It predicts the target model weights for target and keypoint localization ($w_{\text{loc}}$ and $w_{\text{kloc}}$, respectively), along with weights for bounding box and keypoint regression ($w_{\text{br}}$ and $w_{\text{kp}}$, respectively). The transformer architecture, inspired by [48], comprises multiple multi-headed self-attention modules [40], as illustrated in Figure 1(d). Denoting the encoded versions of $f_{\text{test}}$ and $f_{i}$ as $z_{\text{test}}$ and $z_{i}$, respectively, the regression modules take encoded test features as input, along with the target model weights, to predict the final outputs.
Figure 4: Comparison of Mean Squared Error (MSE) and Object Keypoint Similarity (OKS) for individual keypoints against various existing methods: VitPose [47], YOLOv8 [20], TransPose [49], ViPNAS [45], DeepLabCut [21], HRNet [38], and RTMPose [19]. Notably, the marmoset dataset features top-view images of marmosets in a cage with infrequent occurrences of lower limb keypoints. The inherent imbalance in training data leads to notable inaccuracies across all methods for the four bottom-limb keypoints. Moreover, the bottom row displays the output of our approach on a frame from the corresponding dataset’s video sequence, where *-marked frames are from synthetic sequences.
Table 1: Comparison with state-of-the-art methods for pose estimation, including TransPose [49], YOLO-v8 [20], ViPNAS [45], ViTPose [47], DeepLabCut [21], RTMPose [19], and HRNet [38]. We report Mean Squared Error (MSE), Object Keypoint Similarity (OKS), and Probability of Detected Joints (PDJ@x), where $x$ indicates correct keypoint detection if the keypoint lies within a distance of $x\times d$, with $d$ being the diagonal of the bounding box containing the target. We evaluate the models on the APT36K [50], APT10K [53], CrowdPose [24], and DeepLabCut Benchmark datasets [21], which include the TriMouse, Fish, and Marmoset datasets.
Model APT36k Dataset Model APT10k Dataset
MSE OKS PDJ@0.08 PDJ@0.05 MSE OKS PDJ@0.08 PDJ@0.05
TransPose 530.61 0.749 0.747 0.641 TransPose 732.52 0.548 0.621 0.492
YOLOv8-M 459.49 0.798 0.846 0.715 YOLOv8-M 423.89 0.567 0.643 0.544
YOLOv8-L 449.52 0.756 0.856 0.725 YOLOv8-L 406.56 0.567 0.656 0.516
ViPNAS 596.66 0.757 0.729 0.638 ViPNAS 647.12 0.503 0.671 0.503
ViTPose 490.20 0.804 0.916 0.814 ViTPose 548.07 0.562 0.692 0.572
Ours 442.76 0.830 0.936 0.799 Ours 482.69 0.571 0.682 0.565
[…]
Model TriMouse Dataset Model Fish Dataset
MSE OKS PDJ@0.08 PDJ@0.05 MSE OKS PDJ@0.08 PDJ@0.05
TransPose 119.15 0.885 0.925 0.886 TransPose 99.33 0.954 0.913 0.801
YOLOv8-M 82.34 0.901 0.912 0.86 YOLOv8-M 71.15 0.952 0.925 0.823
YOLOv8-L 65.18 0.899 0.928 0.918 YOLOv8-L 59.58 0.952 0.943 0.836
DeepLabCut 102.76 0.897 0.923 0.885 DeepLabCut 92.12 0.958 0.929 0.819
Ours 75.96 0.904 0.946 0.916 Ours 54.24 0.959 0.931 0.842
[…]
Model Marmoset Dataset Model CrowdPose Dataset
MSE OKS PDJ@0.08 PDJ@0.05 MSE OKS PDJ@0.08 PDJ@0.05
TransPose 223.64 0.734 0.796 0.714 RTMPose 241.07 0.715 0.887 0.723
YOLOv8-M 166.31 0.763 0.830 0.744 HRNet 371.41 0.653 0.763 0.625
YOLOv8-L 164.56 0.763 0.839 0.792 ViTPose 339.76 0.681 0.782 0.645
DeepLabCut 190.47 0.768 0.829 0.727 Ours 292.86 0.734 0.861 0.702
Ours 161.95 0.779 0.859 0.784
[…]
3.2.3 Regression Modules
We design the STEP framework to rely solely on the initial bounding box for input dependency. To maintain independence from initial keypoint-related inputs, we introduce the Offset Map Regression Adapter (OMRA) and Gaussian Map Soft Prediction (GMSP) module to incorporate target keypoint embeddings. This allows for simultaneous tracking and pose estimation for the target. We describe each module in detail below:
  • GMSP Module. We utilize a straightforward encoder-decoder architecture, as shown in Figure 2 within the GMSP Module, to learn a soft map relevant to anatomically crucial keypoints of the target. This architecture involves several convolutional and deconvolutional blocks. More specifically, in our experiments, we use two convolutional blocks and two deconvolution blocks for the GMSP Module. As illustrated in Figure 1(b), we train the GMSP module exclusively on training frames, using $\Pi_{kp}$ as supervision to predict a soft map of all possible key points within these frames. This training is performed solely on the training frames, ensuring that the GMSP module is not exposed to test-frame data during the encoding determination process. Given the training features $f_{i}$ encoded with the state encoding of the target state bounding box $\Pi_{b}$, as defined in Equation 4, the transformer predicts model weights $w_{kp}$, which are aware of the target state for the object being tracked. Finally, the adapter module, described below, follows a two-tower architecture to reason the final key points specific to the object of interest.
  • Offset Map Regression Adapter Module. For training frames, the GMSP Module provides information relevant to target state keypoint information to estimate the training feature $f_{i}$. GMSP collaborates with the OMRA Module for the test frame through a two-tower-based architecture, as illustrated in Figure 1(c). In line with [48], the first tower computes an attention map using the target model weights $w_{\text{kp}}$, expressed as $w_{\text{kp}}*z_{\text{test}}$. The second tower handles the output from the GMSP Module. The outputs from both towers are fused using a concatenation operation and then further processed with a CNN. The first tower comprises an attention computation layer followed by three Conv-Norm-ReLU layers. The second tower processes soft predictions from the GMSP Module through four Conv-Norm-ReLU layers. Outputs from both towers are concatenated and then further processed by a CNN with four Conv-Norm-ReLU layers. The final layer predicts the offset map $\hat{K}^{t}_{\text{om}}$. Note that we supervise GMSP Modules solely based on learning from the training frames within the video sequences. While processing subsequent test frames from the video, we freeze all weights associated with the GMSP Module.
    Table 2: Impact on metrics when initializing for target states using a pre-trained detector [55] versus ground-truth boxes.
    Model CrowdPose Dataset
    Initialization
    Per Frame
    Bounding Box
    MSE OKS PDJ@0.08
    STEP Ground Truth 292.86 0.734 0.861
    STEP* Ground Truth 214.87 0.736 0.891
    STEP Detector 292.31 0.731 0.835
    STEP* Detector 213.62 0.733 0.854
    Table 3: Inference speed of our proposed approach compared to existing methods.
    Model Task FPS
    Per Frame
    Detector
    $T_{\text{100}}$ (s)
    STEP Tracking + Pose Estimation 63 1.63
    STEP* Pose Estimation $\sim$63 6.13
    ViTPose [47] Pose Estimation 231 4.97
    TransPose [49] Pose Estimation 43 6.87
    YOLOv8-L [20] Pose Estimation 113 5.43
    YOLOv8-M [20] Pose Estimation 161 5.16
    RTM-Pose [19] Pose Estimation 263 4.92
    DeeplabCut [21] Pose Estimation 68 6.01
    HRNet [38] Pose Estimation 31 7.77
    ViPNAS [45] Pose Estimation 117 5.40
    Figure 5: Frames from supplementary video on Animals [50], Humans [11], and Mouse [21] showcasing tracking and pose estimation. Please refer to the supplementary video for qualitative results on video sequences.
    Table 4: Training on Natural vs Synthetic sequences.
    Model Trained On
    Natural Sequences
    from APT36k
    Synthetic Sequences
    from APT10k
    MSE OKS PDJ@0.08 MSE OKS PDJ@0.08
    Synthetic 374.28 0.635 0.749 395.76 0.764 0.818
    Natural 356.68 0.733 0.872 397.98 0.791 0.861
  • BboxRegressor and Localizer Modules. Localizers and bounding box regressors shown in Figure 1 have a similar architecture. The architectures are depicted in Figure 2 (b), (c), and (d). Initially, they first compute the attention map with their respective target model weights: $w_{\text{br}}$, $w_{\text{loc}}$, and $w_{\text{kloc}}$, which is further processed with three Conv-Norm-ReLU layers. The box localizer $\Lambda_{L}$ predicts localization of target with score map $\hat{B}_{gm}$, the keypoint localizer $\Lambda_{K}$ predicts localization score maps of target keypoints $\hat{K}^{t}_{\textbf{gm}}$, and the Box Regressor $\Lambda_{B}$ predicts dense LTRB offset maps of bounding box extent $\hat{B}_{\text{om}}$.
3.3 Training and Inference
This section outlines the steps involved in training the STEP framework, along with the associated loss functions. We then describe the inference procedure, including the confidence-based memory update, which ensures robust tracking and pose estimation within the STEP framework.
3.3.1 Training
During the training phase, we randomly choose three frames from a video sequence comprising $n$ frames denoted as $\mathcal{S}=\{s_{0},s_{1},\ldots,s_{n}\}$ annotated with bounding boxes and keypoints. Two frames are employed from this sampled set for training, leaving one for testing. For each training frame, we determine target state encodings as defined in Equation 4. The transformer module then predicts the target model weights, which are subsequently utilized by the regressor modules to perform localization, tracking, and keypoint estimation in the test frame. We train the Regressors in the following manner: the $\Lambda_{O}(.)$ predicts $\hat{K}^{t}_{\text{om}}$, the $\Lambda_{K}(.)$ predicts $\hat{K}^{t}_{\text{gm}}$, the $\Lambda_{B}(.)$ predicts $\hat{B}_{\text{om}}$, and the $\Lambda_{L}(.)$ predicts $\hat{B}_{\text{gm}}$ for the test frame. The respective ground truths for these predictions are $K^{t}_{\text{gm}},K^{t}_{\text{om}},B_{\text{om}}$, and $B_{\text{gm}}$ obtained as discussed in Section 3.2. For $B_{\text{gm}}$ and $K_{\text{gm}}$, we employ target classification as described in [2]. We use a generalized Intersection over Union loss to predict $B_{\text{om}}$. Additionally, we apply mean-squared error loss to penalize the keypoint regression module for predicting $K^{t}_{\text{om}}$. Based on training-frame information, the GMSP Module is expected to effectively infer an enhanced target state of keypoints for the test frame. Therefore, we train the GMSP Module only on training frames to predict $\hat{K}_{\text{gmsp}}$ and freeze its weight while processing the test frame. It’s crucial to note that the loss applied to the regressors is derived explicitly from predictions made for the test frame, while the loss for the GMSP Module is based on predictions from the train frames. We now provide a detailed description of the loss functions for training in Section 3.3.2, which supervise the estimation of target-aware model localization weights $w_{loc},w_{kloc}$ and offset map weights $w_{br},w_{kp}$, ensuring the end-to-end optimization of all learnable parameters in the model.
3.3.2 Loss Functions
We utilize the target classification loss, denoted as $\mathcal{L}_{\text{cls}}$ similar to [2] for the Gaussian map predicted in the test frame through bounding box and keypoint regressors. The GMSP module is trained using the loss $\mathcal{L}_{\text{cls}}$, and this training is conducted on the training frames $\mathcal{S}_{\text{train}}$ to predict map $\hat{K}_{\text{gmsp}}$, as opposed to the test frames. For the regression of offset maps associated with keypoints, we employ a hinged mean-squared loss denoted as $\mathcal{L}_{\text{hom}}$. This loss is only applied when the keypoint is visible in the annotated ground truth. We employ generalized Intersection over Union loss [35] using LTRB bounding box representation for bounding box offset map regression. Loss functions employed during the training phase of STEP are defined in equation 7. In our experiments, we use $\lambda_{1}=\lambda_{3}=\lambda_{5}=100$ and $\lambda_{2}=\lambda_{4}=10$. To generate ground-truth offset maps for $K_{\text{om}}$ and $B_{om}$, we use a stride value $s=16$ and Gaussian Maps $B_{\text{gm}}$ and $K_{\text{gm}}$ are generated at a spatial resolution of $18\times 18$.
$\mathcal{L}_{\text{test-frame}}$ $=\lambda_{1}\mathcal{L}_{\text{cls}}(B_{\text{gm}},\hat{B}_{\text{gm}})+\lambda_{2}\mathcal{L}_{\text{giou}}(B_{\text{om}},\hat{B}_{\text{om}})$
$+\lambda_{3}\mathcal{L}_{\text{cls}}(K^{t}_{\text{gm}},\hat{K}^{t}_{\text{gm}})+\lambda_{4}\mathcal{L}_{\text{hom}}(K^{t}_{\text{om}},\hat{K}^{t}_{\text{om}})$ (5)
$\mathcal{L}_{\text{train-frame}}=\mathcal{L}_{\text{gmsp}}=\lambda_{5}\mathcal{L}_{\text{cls}}(K_{\text{gm}},\hat{K}_{\text{gmsp}})$ (6)
$\mathcal{L}_{\text{tot}}=\mathcal{L}_{\text{test-frame}}+\mathcal{L}_{\text{train-frame}}$ (7)
3.3.3 Inference
During the inference phase, our approach requires initialization with a bounding box around the target object for the initial frame $s_{0}\in\mathcal{S}$ to establish target state information. This initialization could be done manually or with existing class-specific detectors. For the first frame, our training set is prepared by duplicating $s_{0}$ such that $\mathcal{S}_{\text{train}}=\{s_{0},s_{0}\}$. Subsequently, our network predicts the sequence’s next frame $s_{1}$ output. Following, this prediction $s_{1}$ along with the predicted target state encodings, modifies our network’s memory for training, forming $\mathcal{S}_{\text{train}}=\{s_{0},s_{1}\}$. We modify $\mathcal{S}_{\text{train}}$ according to confidence thresholding, specifically when $n$ keypoints are predicted with confidence exceeding a certain threshold $\tau_{m}$ in the map $\hat{K}^{t}_{\text{gm}}$. When there’s no memory update, the test frame is solely processed through the transformer, which is sufficient, enhancing performance speed. This is because the training frames and their corresponding target encodings remain unaltered. This iterative process continues until the end of the sequence. To determine the exact coordinates of keypoints and the bounding box in image space from the regressed offset maps $\hat{K}^{t}_{\text{om}}$ and $\hat{B}_{\text{om}}$, we first extract the center location from the predicted confidence maps $\hat{K}^{t}_{\text{gm}}$ and $\hat{B}_{\text{gm}}$ by identifying the coordinate with the maximum value. We then query the offsets regressed at that location in $\hat{K}^{t}_{\text{om}}$ and $\hat{B}_{\text{om}}$, map it back into the image space of $s_{i}\in\mathrm{R}^{H_{i}\times W_{i}\times 3}$ by taking downsampling factor in Equation 2 into account and obtain keypoints and the bounding box.
Implementation details. We employ Adam optimizer with an initial learning rate set to $10^{-4}$, decayed by a factor of $0.1$ after 50 epochs. We update the memory state of our network during inference when at least $50\%$ of the total keypoints and the target localization both are predicted with a confidence value greater than $\tau_{m}=0.6$. All experiments were performed using NVIDIA GeForce RTX-4090 GPUs.
Table 5: Impact of the GMSP Module and $\phi_{\text{kp}}$.
Model Configuration $\phi_{\text{kp}}$ Keypoint Initialization APT36k Validation Split
MSE OKS PDJ@0.08
w GMSP 356.68 0.733 0.872
w/o GMSP 464.78 0.698 0.832
w GMSP 372.94 0.714 0.841
w/o GMSP 468.78 0.628 0.802
Table 6: Impact of inclusion of embeddings in training features $f_{i}$
$\phi_{\text{loc}}$ $\phi_{\text{test}}$ $X_{\text{test}}$ $\phi_{\text{kp}}$ MSE OKS PDJ@0.08
356.68 0.733 0.872
393.54 0.686 0.765
352.82 0.724 0.833
396.73 0.682 0.763
364.2 0.701 0.785
Figure 6: We demonstrate tracking and pose estimation during the occlusion of a target human instance by another human with a similar appearance. Zoom-in is recommended for clarity. Panel (a) illustrates memory updates occurring only when at least 50% of keypoints and the bounding box localization are predicted with a probability greater than $\tau_{m}=0.6$. In contrast, panel (b) shows memory updates being performed regardless of the probability value. In both panels, frames with low confidence, which may introduce or propagate errors, are indicated by a red bounding box. In panel (b), STEP wrongly adjusts its confidence from the original target to the new instance based on the memory state. Subsequently, it starts tracking and performing pose estimation on the new object.
Table 7: Impact of the number of training frames used, $|\mathcal{S_{\text{train}}}|.$
#Training frames APT36k (Val Split) APT10k (Val Split)
MSE OKS MSE OKS
1 initial 465.48 0.697 536.69 0.684
1 initial + 1 recent 360.87 0.723 397.30 0.768
1 initial + 2 recent 358.31 0.723 395.82 0.778
Rolling recent-2 495.89 0.682 543.49 0.673
Rolling recent-3 492.62 0.674 538.95 0.686
Conf. Rolling-2 356.68 0.733 395.76 0.764
Conf. Rolling-3 357.38 0.742 394.89 0.764
4 Results and Discussion
We assess our model across multiple datasets encompassing various species. APT36K [50], consisting of 2,400 video clips collected and filtered from 30 animal species with 15 frames for each video, resulting in 36,000 frames in total for animal pose estimation and tracking, whereas APT10K [53] consists of 10,015 images collected and filtered from 23 animal families and 60 species. Our evaluation also spans across DeepLabCut Benchmark Datasets [10] [21] of diverse species in the lab setting, which includes the Tri-mouse dataset with 161 labeled frames where three wild-type male mice ran on a paper spool following odor trails; the Marmosets dataset with 7,600 labeled frames from 40 different marmosets collected from 3 different colonies, and Fish dataset with 100 labeled frames of Schools of inland silversides (Menidia beryllina). We utilized the CrowdPose [24] dataset, containing about 20,000 images and a total of 80,000 human poses with 14 labeled key points to assess the performance of STEP in human-based pose estimation and tracking.
4.1 Data Processing
Our network is trained by sampling frames from sequences in datasets, as detailed in Section 3.3. We generate sequences from image-based datasets like APT10K and CrowdPose by applying random affine transformations during training. We create smooth transitions with random affine transformations for validation splits, resulting in 15-frame video sequences. These sequences are saved to evaluate the performance of existing methods. Figure 3 visually showcases qualitative tracking and pose estimation results on a sequence from the APT36k Dataset. The validation splits are formatted as $(n_{v},n_{tid},n_{f})$, representing the number of video sequences, unique targets across all videos, and frames per video, respectively. Across the datasets— APT36k, APT0k, CrowdPose, Fish, TriMouse, and Marmoset— these splits contain $(237,348,15)$, $(201,269,15)$, $(1572,6803,15)$, $(24,336,15)$, $(24,78,15)$, and $(199,239,15)$ data configurations, respectively. We train distinct models for each dataset to conduct both quantitative and qualitative evaluations. Since the different datasets utilize different keypoint landmarking systems, adjustments are necessary for the final layers of the Regressors in STEP. These minor modifications enable efficient fine-tuning and facilitate the model’s rapid adaptation to new datasets. However, we present results based on training each model from scratch for fair comparison. Our approach successfully learns to predict poses for different species within a dataset (e.g., APT36K, which includes 30 species) as they adhere to a consistent set of keypoint landmarks. All methods listed in Table 1 also follow a similar evaluation setup, except for ViTPose. ViTPose employs joint training across all datasets to train a shared encoder and then employs dataset-specific decoders to regress keypoints.
4.2 Quantitative Evaluation.
In Table 1, we present a comparison of STEP with state-of-the-art methods on pose estimation in terms of Mean Squared Error (MSE), Object Keypoint Similarity (OKS), and Percentage of Detected Joints (PDJ). PDJ@$x$ signifies a correct keypoint detection if it lies within the $x\times d$ distance, where $d$ is the diagonal of the bounding box containing the target. The top two performances are highlighted in bold. In Ours, we use only the initial frame’s bounding-box information and perform simultaneous tracking and pose estimation. In Ours*, we use per-frame ground-truth bounding box annotations to scale, crop, and center the images appropriately, similar to existing top-down approaches, to mitigate tracking errors and ensure a fair comparison of pose metrics. When evaluated using a top-down approach, our approach (Ours*) outperforms existing state-of-the-art methods and demonstrates competitive performance in pose estimation while simultaneously performing the tracking task also (Ours), as shown in Table 1. Complementing the information in Table 1, Figure 4 displays the evaluation of Mean Squared Error (MSE) and Object Keypoint Similarity (OKS) scores for every individual keypoint. In Figure 8(b), a comparative assessment of tracking results is provided on the APT36K Dataset with existing tracking methods. The evaluation involves Intersection Over Union (IoU) and Overlap Precision (OP) of bounding boxes with thresholds ranging from $(0,1]$ between STEP and existing tracking methods.
4.2.1 Automatic Initialization
Our proposed approach requires initialization with a bounding box to establish information on the target of interest. We conduct an experiment to assess the effect of initialization by employing a pre-trained Detector [55] to identify humans in the validation split of the CrowdPose Dataset [24]. The detector provides class-specific bounding boxes, and since CrowdPose is a human dataset, we filter out the predicted bounding boxes specific to the person category. However, the detector can sometimes mistakenly identify non-person objects as human or detect instances whose corresponding keypoint annotations are unavailable in the ground truth, which can hinder metric evaluation. To address this, we filter the predicted bounding boxes by selecting those with the maximum Intersection over Union (IoU) with the annotated boxes in the dataset. We run STEP on validation splits with initializations from the detector and with annotation available in the CrowdPose dataset and observe the minimal impact on metrics, as evident from Table 2. Initializing STEP with a pre-trained detector [55] has minimal impact on performance while fully automating the process.
4.3 Qualitative Results
A supplementary video is available, demonstrating the results on video sequences of animals from [50], humans from [11], and mice and fishes from [21]. We present a frame from the sequences in the supplementary video in Figure 5.
4.4 Inference Speed Analysis
For inference speed, We evaluate all methods on an equivalent number of frames from the respective datasets, matching the number of frames in the validation split of APT-36K, as shown in Table 3. Given the differences in evaluation settings for pose estimation across SOTA methods, we established the following protocols to asses comparison: we fix the resolution of all frames to $(H\times W=256\times W)$ maintaining the aspect ratio for STEP (Ours). For Deeplabcut and STEP* (Ours*), we cropped detected instances into square patches and resized them to $(256\times 256)$. For the other methods listed in Table 1, we employed cropping similar to that used in ViTPose, i.e., more horizontal cropping for animals $(192\times 256)$ and more vertical cropping for humans $(256\times 192)$. Note that our proposed method, STEP, achieves an average speed of 63 FPS for both tracking and pose estimation tasks while requiring initialization for the first-most frame only. In contrast, STEP* and other top-down approaches face a performance bottleneck of per-frame detections. For object detection, we employ well-established detectors, specifically MegaDetector [15] for animal detection and Dino [55] for human detection, achieving inference speeds of $29$ FPS and $22$ FPS, respectively. In Table 3, We include $T_{100}$ to clearly demonstrate the total time required to process $100$ frames when both tasks are performed sequentially with Dino Detector, facilitating a more straightforward assessment. Furthermore, we would like to highlight that both STEP and STEP* maintain tracking information across videos. In contrast, other methods necessitate an additional step to establish target correspondence across frames, such as tracklet stitching in DeepLabCut with an FPS of 2400 [21]. FPS evaluations were conducted with a batch size of 1, and all experiments were performed on a device equipped with an NVIDIA Geforce RTX-4090.
The video also contains results using the top-down approach with STEP*, utilizing per-frame bounding box information. Since the detections for each frame are known in advance, STEP* can manage newly entered objects, maintain tracking, and estimate their respective poses.
4.5 A Case Study on AMC Live Streams
Utilizing STEP, we extracted self-motion and pose information from live video streams at the Awajishima Monkey Center (AMC) in Hyogo prefecture, Japan (Video recordings used with permission, Live Stream: Link). The center accommodates a large group of freely moving macaque monkeys, capturing their behavior through a live camera overlooking a popular visitor area. While macaque activity is extensively studied in animal ethology, typically at a coarse temporal scale, focusing on proportions of time spent on various activities like foraging and grooming [17], our emphasis was on a finer measurement of self-motion. We aimed to leverage STEP to understand alterations in monkey behavior throughout the day and in response to events such as feeding. Observations reveal that monkeys generally move at median speeds below $1$ m/sec (Figure 2(a)), but they can also exhibit rapid movements (Figure 2(b)). Elevated activity is noticeable during feeding episodes, as depicted in Figure 2(c). Figure 8(a) illustrates the movement speed observed across different track-ids on streams over two days. It’s important to note that the analysis is confined to 2D-image space, excluding consideration of the third dimension. We employed an existing detector [29] (fine-tuned on monkeys) to initialize the STEP tracking and estimate the pose for subsequent frames. STEP demonstrates satisfactory performance, effectively quantifying behavioral changes in a setting with high variability in animal occupancy, scale, and pose.
Figure 7: (a) slowest and (b) fastest quartiles of self-movement velocities over recordings from 2 days. The provision of feeding results in animals moving towards food, causing movement from black to white clusters, illustrated in (c). Trajectories are plotted by tracking the center of the bounding box across all selected tracks. A high occupancy representative scene with overlaid tracks and poses is shown in (d).22footnotemark: 2
Figure 8: (a) Movement velocity observed overall recorded data from AMC (tracks longer than 1 sec). Increased movement typically aligns with feeding episodes. Distances are derived from AMC’s perspective on Google Maps. (b) IoU and Overlap Precision $(\text{OP}_{T})$ for bounding box tracking in STEP compared with existing methods: DiMP [2], ToMP [31], ATOM [8], and TAMOS [32].
4.6 Ablation Study
We conducted multiple ablation studies to assess the performance of the proposed STEP approach, utilizing the APT36k and APT10k datasets. The reported metrics showcase the results obtained when all variants were trained for 100 epochs and subsequently evaluated on a reduced validation split consisting of 150 video sequences from each dataset.
4.6.1 Natural vs Synthetic Data sequence.
Table 4 presents metrics illustrating the impact on performance when training the model on natural and synthetic sequences. The model effectively generalizes natural video sequences even when trained on synthetic sequences.
4.6.2 Architectural Ablations
In Table 5, we examine the influence of the GMSP module. We hypothesize that the GMSP module learns soft predictions for the keypoints. To validate this, we experimented by removing the GMSP module and replacing it with ground-truth keypoints. In this configuration, we omitted the branch injecting features from GMSP to regression modules for test frames. This setup resulted in a decrease in overall performance. The inclusion of the GMSP module not only eliminates the need for keypoint initialization but also contributes to increased performance. Additionally, the soft predictions generated by GMSP provide more enriched information for keypoint encodings than the keypoint initialization setting. The incorporation of $\phi_{\text{loc}}$ has been demonstrated to improve performance, as indicated in [31]. Similarly, we explore the impact of including $\phi_{\text{kp}}$ in determining target state encodings. As presented in Table 5, a notable performance enhancement in pose estimation is attributed to incorporating the learnable embedding for each keypoint. Furthermore, in Table 6, we provide a more detailed assessment of the impact of test-frame embedding $\phi_{test}$, bounding box embedding $\phi_{loc}$ and on the inclusion of test-frame $X_{test}$ for model weight prediction.
4.6.3 Memory Update
In Table 7, we examine the impact of the number of training frames for target model weight predictions during inference. By updating training frames $\mathcal{S}_{\text{train}}$ using confidence thresholding $\tau_{m}$ as detailed in Section 3.3, we observe a performance improvement. In Figure 6, we demonstrate the robustness of memory update for accurate tracking and pose estimation of target objects.
4.7 Limitations
Our proposed approach requires initialization through a bounding box for the first-most frame in a given video to infer target state encodings. This initialization establishes the target of interest and could be manual or fully automated using object detectors such as [29] [55]. We observed that the network tends to estimate the pose keypoints within the predicted bounding box or, conversely, predicts a bounding box that encompasses the predicted keypoints. As our approach operates simultaneously, errors originating from one task can negatively impact the accuracy of the other task. This is evident from the metrics evaluated between ”Ours” and ”Ours*” in Table 1.
5 Conclusion
This work proposes a novel framework for Simultaneous Tracking and Estimation of Pose across diverse animal species and humans. Our method estimates the pose of the object of interest while simultaneously tracking it, utilizing a Transformer-based discriminative target model prediction approach. The key contributions of our framework include the Gaussian Soft Map Prediction module and the Offset Map Regression Adapter module, which effectively removes the dependency on keypoint initialization to estimate target state encodings. We show that the network equipped with the proposed modules performs better than the version relying on keypoint initialization. Unlike traditional top-down approaches that depend on per-frame detections for pose estimation and are thus limited by speed bottlenecks, our method overcomes this issue by relying on efficient tracking. The proposed STEP framework achieves state-of-the-art performance in both tracking and pose estimation tasks. Lastly, we showcase the robustness of our approach in challenging video scenarios with high animal occupancy, underscoring its effectiveness in both tracking and pose estimation.

### v2

@@
 These modules remove the necessity of keypoint target states as input, streamlining the process.
-Our method starts with a known target state initialized through a pre-trained detector or manual initialization in the initial frame of a given video sequence.
+Our method starts with a known target state in the initial frame of a given video sequence.
 It then seamlessly tracks the target and estimates keypoints of anatomical importance as output for subsequent frames.
@@
 In summary, our contributions are as given below.
-  • We conduct a comprehensive study on the pose estimation task within a discriminative model prediction framework.
+  • We conduct a comprehensive study on pose estimation task within a discriminative model prediction framework.
 We utilize a Transformer-based model prediction module while simultaneously tracking the object of interest— an unexplored area in previous research.
@@
 A failure case is showcased in the last frame of (c) for the red bounded target, stemming from ambiguous activations within the red target’s score map, as evident in the last frame of (a).
-Additional qualitative results across various species are provided in the Supplementary material.
 3.2 Network Architecture
@@
 Initializing STEP with a pre-trained detector [56] has minimal impact on performance while fully automating the process.
+Furthermore, our work addresses the fundamental challenge of simultaneous tracking and pose estimation in wild settings.
+We highlight that in constrained environments, such as controlled laboratories or zoo enclosures, animal localization is relatively straightforward, as most or all elements in the scene remain static.
+Simple bottom-up approaches can efficiently handle pose estimation and localization tasks [17], which can be easily integrated into our method and can be utilized as an initialization step.
 4.3 Qualitative Results
-A supplementary video is available, demonstrating the results on video sequences of animals from [50], humans from [11], and mice and fishes from [21].
+A supplementary video is available (Link) demonstrating the results on video sequences of animals from [51], humans from [11], and mice and fishes from [22].
 We present a frame from the sequences in the supplementary video in Figure 5.
+The video also contains results using the top-down approach with STEP*, utilizing per-frame bounding box information.
+Since the detections for each frame are known in advance, STEP* can manage newly entered objects, maintain tracking, and estimate their respective poses.
 4.4 Inference Speed Analysis
@@
 FPS evaluations were conducted with a batch size of 1, and all experiments were performed on a device equipped with an NVIDIA Geforce RTX-4090.
-The video also contains results using the top-down approach with STEP*, utilizing per-frame bounding box information.
-Since the detections for each frame are known in advance, STEP* can manage newly entered objects, maintain tracking, and estimate their respective poses.
 4.5 A Case Study on AMC Live Streams
@@
 Lastly, we showcase the robustness of our approach in challenging video scenarios with high animal occupancy, underscoring its effectiveness in both tracking and pose estimation.
+Acknowledgments
+We thank Prof.
+Kazunori Yamada from Osaka University and Hisami-san from the Awaji Island Monkey Center for granting us access to the center’s video feeds for our analysis.
+We also thank Dr.
+Mami Matsuda for facilitating communication with the Awaji Monkey Center.
+Additionally, we acknowledge the support of the Jibaben Patel Chair in AI for this work.
+Data Availability
+All experiments are performed on publically available datasets: APT36K [51] (https://github.com/pandorgan/APT-36K), APT10K [54] (https://github.com/AlexTheBad/AP-10K), CrowdPose [25] (https://github.com/jeffffffli/CrowdPose) and TriMouse, Marmoset and Fish from DeepLabCut benchmark dataset [22] https://benchmark.deeplabcut.org/datasets.html.
+For Figure 2 and Figure 8(a), we used the video stream available at (https://www.youtube.com/live/lsxYH2XQQCg).
+We have made the STEP project page publicly accessible at Link, which includes access to supplementary video (https://youtu.be/TxSk3OtJ6R8) and code (https://github.com/shash29-dev/STEP).